OpenStudy (anonymous):

Solve the equation by factoring: x^2 + 5x ­ - 24 = 0 Is the answer -3, 8 or -8, 3???

OpenStudy (photon336):

You can do this two ways by factoring or by using the quadratic formula \[(x+8)(x-3) = 0\]

OpenStudy (photon336):

there are two solutions to this equation @jojor_124 can you see this?

OpenStudy (photon336):

solving this is easy \[\frac{ (x+8) }{ (x+8) }*(x-3) = \frac{ 0 }{ (x+8) }\] we divide both sides by (x+8) and we're left with this. how would we solve this? \[(x-3) = 0 \] and \[(x+8) = 0 \]